a. Show that a given change in the money stock has a larger effect on output the less interest-sensitive is the demand for money. Use the formal analysis of Section 11-5.
b. How does the response of the interest rate to a change in the money stock depend on the interest sensitivity of money demand?
1
Expert's answer
2021-08-24T14:05:04-0400

a)

"L=kY-hi" (demand equation)

It shows that h is interest sensitive.

"Y=r\bar{A}+\frac{b}{h}\frac{\bar{M}}{p}\\Y=r×\frac{b}{m}" (Equilibrium output in is-lm model)

It shows that given the money stock,the less interest sensitive demand for money lie,lower h increases changes the value of output.


b)

Equation of money in the market is given as

"I=\frac{I}{h}[kY-\frac{\bar{M}}{P}]\\i=\frac{I}{h}×kY-\frac{I}{h}×(\frac{\bar{M}}{P})\\\frac{i}{m}=-\frac{I}{h}"

As shown by the equition,a change in interest rate occurs due to change in output which depends on interest sensitivity of Money demand (h).